A nationwide survey found that 64% of people in the Kingdom like Berger. If 3 people are selected at random with replacement, what is the probability that all three don't like Berger?
Answer #1

P(like berger) = 0.64

P(don't like berger) = 1 - 0.64 = 0.36

So, P(all 3 don't like berger) = 0.36*0.36*0.36 = 0.046656

I hope this will help. Thank you.